OpenStudy (anonymous):

Help If sin(x) = 1/3 and sec(y) = 25/24 , where x and y lie between 0 and π/2, evaluate the expression using trigonometric identities. (Enter an exact answer.) What I know: cos (y) = 25/24 sin (x + y) = (sinx)(cosy) + (cosx)(siny) so... sin (x + y) = (1/3)(24/25) + (cosx)(siny) How do you get cosx and siny? Thanks.

hartnn (hartnn):

u have sin x use this to get cos x \(\huge sin^2 x+cos^2x=1\)

hartnn (hartnn):

similarly u have cos y , find sin y using this : \(\huge sin^2 y+cos^2y=1\) can u ?

hartnn (hartnn):

@franciscanmonk did u understand that ^^ ?

OpenStudy (anonymous):

hmm i see. okay ill use it

OpenStudy (anonymous):

so... if i use that identity... i'd get cosx = sqrt(8/9) siny = (7/25) so.. sin(x+y) = (1/3)(24/25) + (sqrt(8/9))(7/25) sin (x+y) = so... final answer is (24-7sqrt(8))/25 right?

OpenStudy (anonymous):

sorry. (24-7sqrt(8))/75

hartnn (hartnn):

something's wrong.....why minus in between ?

OpenStudy (anonymous):

oh sorry its + haha

OpenStudy (anonymous):

hmmm... OH i see. sqrt(8/9) is (2sqrt(2)/3) (8/25) + (7/25)(2sqrt(2)/3) (8/25) + (14sqrt(2)/75) (24/25) + 14sqrtr(2)/75 (24 + 14 sqrt(2))/75) ....right?

hartnn (hartnn):

thats perfect !

OpenStudy (anonymous):

AWESOME!! Thank you!!!!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

haha forgot sin^2x + scos^2x = 1